OpenStudy (anonymous):

Last one ... :-) An equation of a circle is x2 + y2 + 10x - 6y + 18 = 0. Show all your work in determining the center and radius of this circle. In complete sentences, explain the procedure used.

OpenStudy (ghazi):

whilst determining center 2*g*x= 10x and 2*f*y=-6y now calculate g and f after that use r= square root of (h^2+k^2)

OpenStudy (anonymous):

still confused can you break it down a little more for me ?

OpenStudy (ghazi):

did you solve to find center of circle i mean did you find g and f

OpenStudy (anonymous):

how do you solve for g and f?

OpenStudy (anonymous):

could you please explain it to me?

OpenStudy (anonymous):

Move 18 to the RHS and complete the square for the x and y expressions.\[x^2+10 x+25+y^2-6 y+9=-18+25+9 \]\[(x+5)^2+(y-3)^2=16 \](-5,3) is the circle's center and the radius is radical 16 or 4. A plot is attached.

OpenStudy (anonymous):

You're the best Thanks

OpenStudy (anonymous):

You're welcome.